Vector Mechanics Dynamics

The 6.85 lb disk A has a radius of 4.35 in and is idle. The 12.09 lb disk B has a radius of 7.49 in and an angular velocity of 910 when it contacts the disk A. Neglecting the friction in the bearings, determine the total thrust "Ft" exerted by the friction force on the disco A.

Р A Фо ТА ТВ
Answer #1

Feel free to ask any question but remember to rate the answer.

Moment of Inertia of dist A IS 2 IA= L x (WA XA2 g Th 5 ( -X(6.85X4.2 12 IA= 0.01397 lb. A g2 32.2 M. I of disk B IS, IB 1X (

Moment about point A IS ot (Ft) VA - TAWA ft= 0.01397.WA. (4.356127 Ft= 0.03855WA FBD of disk B is as follows. IB WD fowe +Nt

Taking moment about point B, Тв. 00 — (ft). Y() = Тg bg — - Ig. oo –Co. 039) он): p = Трор (o od313) (35. 2999) — (o-o381 X7:

50-5y07 од = 9 - 9946 — 0:3190 он *О. 9o92 0A = 9. 1946 : LDA= 104.7453 radly — Toy-453x ( 6 ) • w = [000. Дуо тут But, WB =